$SPAD/src/input multiple.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool multiple.output
)set message test on
)set message auto off
)clear all
--S 1 of 8
draw(sin(x),x=0..2*%pi)
--R
--R   Compiling function %B with type DoubleFloat -> DoubleFloat
--R   Graph data being transmitted to the viewport manager...
--R   AXIOM2D data being transmitted to the viewport manager...
--R
--R   (1)  TwoDimensionalViewport: "sin(x)"
--R                                                 Type: TwoDimensionalViewport
--E 1
--S 2 of 8
v1 := %
--R
--R
--R   (2)  TwoDimensionalViewport: "sin(x)"
--R                                                 Type: TwoDimensionalViewport
--E 2
--S 3 of 8
draw(cos(x),x=0..2*%pi,curveColor==light red())
--R
--I   Compiling function %D with type DoubleFloat -> DoubleFloat
--R   Graph data being transmitted to the viewport manager...
--R   AXIOM2D data being transmitted to the viewport manager...
--R
--R   (3)  TwoDimensionalViewport: "cos(x)"
--R                                                 Type: TwoDimensionalViewport
--E 3
--S 4 of 8
v2 := %
--R
--R
--R   (4)  TwoDimensionalViewport: "cos(x)"

--R                                                 Type: TwoDimensionalViewport
--E 4
--S 5 of 8
graphs v1
--R
--R
--R   (5)
--R   [Graph with 1 point list, undefined, undefined, undefined, undefined,
--R    undefined, undefined, undefined, undefined]
--R                                    Type: Vector(Union(GraphImage,undefined))
--E 5
--g1 := elt(graphs v1,1)::GraphImage
--S 6 of 8
g1 := getGraph(v1,1)
--R
--R
--R   (6)  Graph with 1 point list
--R                                                             Type: GraphImage
--E 6
--S 7 of 8
putGraph(v2,g1,2)
--R
--R                                                                   Type: Void
--E 7
--S 8 of 8
makeViewport2D(v2)
--R
--R   AXIOM2D data being transmitted to the viewport manager...
--R
--R   (8)  TwoDimensionalViewport: "cos(x)"
--R                                                 Type: TwoDimensionalViewport
--E 8
)spool
)lisp (bye)

———-

References^{[1] nothing}

5